OpenStudy (anonymous):

Let f(x,y)= 5/ (x^2 + y^2). Find an equation of a linear approximation to f(x,y) at (-1,2,1) and use it to estimate f(-1.05,2.1)

OpenStudy (amistre64):

wouldnt a linear approx just be a plane?

OpenStudy (amistre64):

in fact, the plane is worked up similar to that of pont slope formula; except its point slope slope formula :)

OpenStudy (amistre64):

\[z-z_o=f_x(x-x_o)+f_y(y-y_o)\]

OpenStudy (anonymous):

not sure @amistre64!

OpenStudy (anonymous):

which one's x and whichone's x not etc.

OpenStudy (amistre64):

\[f(x,y)=\frac{5}{x^2+y^2}\] \[f_x=\frac{-5(2x)}{(x^2+y^2)^2}\] \[f_y=\frac{-5(2y)}{(x^2+y^2)^2}\]

OpenStudy (anonymous):

i'm sorry! can you possibly explain how you got fx and fy??

OpenStudy (amistre64):

ok, let me post this and ill try given f( x=-1.05,y=2.1) that gives us: \[f_x=\frac{10(1.05)}{(1.05)^2+(2.1)^2}\] \[f_y=\frac{-10(2.1)}{(1.05)^2+(2.1)^2}\]

OpenStudy (amistre64):

do you know partial derivatives?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

fx means we treat y as a constant \[\frac{5}{x^2+y^2}=5(x^2+c)^{-1}\] and derive it y is the same except for make x^2 = c

OpenStudy (anonymous):

oh ok!

OpenStudy (amistre64):

is that an OHHH ok? or one of those, i spose ... oks ?

OpenStudy (amistre64):

in essense we ignore the variable that is not important in the partial and treat it as any other constant

OpenStudy (amistre64):

they give us x0 and y0; to find z0 we solve f(x,y) for the given point and call f(x,y) = z

OpenStudy (anonymous):

dat wuz an i kinda get it ok

OpenStudy (amistre64):

if your confident with your usual derivative stuff; its just a little extra thinking to ignore that variable

OpenStudy (amistre64):

f(x,y) = xy^2 fx = y^2 ; since Cx derives to C and C=y^2 Fy = 2xy since Cy^2 = 2Cy and C = x in that case

OpenStudy (anonymous):

i get it!

OpenStudy (amistre64):

lol, good :)

OpenStudy (amistre64):

i knew this looked familiar

OpenStudy (anonymous):

@amistre64 LOL Yup! i came in not that long ago, and saw the bump button available, so i pushed it! LOL

OpenStudy (amistre64):

good thing it wasnt shiny and red :)

OpenStudy (anonymous):

LOL:)